# Mixed-wet percolation on a dual square lattice

### v1

@@ 1 Introduction
Percolation theory was first recognized as a critical phenomenon by Broadbent and Hammersley in 1957 [1]. Little did they know that their paper would give rise to literally thousands of papers in the ensuing years. It remains the foremost example of a non-thermal critical phenomenon [2, 3]. The percolation problem in its primal form is easy to grasp. However, the percolation problem comes in many disguises. We may list some. Ambogaekar, Halperin and Langer demonstrated that the conductivity of strongly disordered conductors is related to a percolation threshold [4]. A similar line of reasoning for measuring the permeability of porous rocks is behind the well-known Katz and Thompson method of mercury injection [5]. Slow injection of one fluid into a porous medium already containing another fluid, and the two fluids are immiscible, is well modeled by the invasion percolation model [6]. A more recent version of the percolation problem that has generated considerable interest is explosive percolation [7].
@@
We present here a new percolation model that is motivated by the flow of immiscible fluids in a mixed-wet porous medium. Percolation models have widely been used in the past and present to characterise porous media flow, especially in the capillary-dominated regime where the flow is governed by the disorder in the pores [8, 9, 10, 11, 12]. For example, the displacement of an immiscible fluid by another fluid inside a porous medium during capillary-dominated flow produces fingers due to capillary instabilities. These fingers are called capillary fingers, which were characterized by the invasion percolation model [13, 14]. The invasion percolation model has also been used to find the relative permeability and saturation profile near the percolation critical point for a static two-phase flow system in the capillary-dominated regime [15]. The percolation model we present here is motivated by the experimental studies [16, 17], and the computational modeling [18] of two-phase flow in an engineered mixed-wet porous medium that is made of a random mixture of two types grains with opposite wetting properties with respect to the fluids flowing in the medium. Imagine a random mixture of two types of grains, say $A$ and $B$, and two immiscible fluids, say fluid $1$ and fluid $2$. Suppose the grain type $A$ is less wetting to fluid $1$ compared to fluid $2$, i.e., the wetting angle $\theta_{2A}<\pi/2$, see Fig. 1. The grain type $B$, on the other hand, is more wetting to fluid $1$ compared to fluid $2$, i.e., $\theta_{1B}<\pi/2$. Suppose now we have an interface between fluid $1$ and $2$ moving in a pore with $A$-type walls. The capillary force between the two fluids will then point towards fluid $1$ as fluid $2$ is more wetting than fluid $1$, see Fig. 1 (left). If the walls are both of type $B$, the capillary force at the interface will point towards fluid $2$. However, if one wall is of type $A$ and the other of type $B$, we have the situation shown in Fig. 1 (right). In this case, the capillary forces will be much weaker than the two previous cases as the contribution from each wall essentially cancels.
@@
Geistlinger et al. [19] did a micro-computed tomography ($\mu$-CT) study of a mixture of sand grains with contact angle either being $0^{\circ}$ or $100^{\circ}$ with respect to the fluid interfaces. They then injected one of the fluids into the ensuing porous medium saturated with the other fluid. They would recognize a structural percolation transition in the clusters of the defending fluid after the invasion process. Irannezahad et al. [16, 17] studied the same problem numerically but not in a percolation context. Fyhn et al. [18] studied steady-state flow in mixed-wet porous media using a dynamic pore-network model [20]. By steady-state flow, we mean flow where the macroscopic averages of the various variables are constant or fluctuate around well-defined averages, and there are no macroscopic saturation gradients present [21]. Fyhn et al. [18] recognized that the steady-state mixed wet problem is a percolation problem in disguise.
@@
In 2009, Tallakstad et al. [22, 23] demonstrated that there is a flow regime under immiscible steady-state two-phase flow in porous media where the flow rate is proportional to the pressure gradient raised to a power close to two. This is due to interfaces held in place by capillary forces would be mobilized with increasing pressure drop. Sinha and Hansen [24] found that if pressure drop rather than flow rate was the control variable, there typically would be a threshold pressure drop for flow to occur. This would be caused by the immobilized interfaces blocking the flow. The presence of this threshold would make the determination of the exponent in the power law between flow rate and pressure drop difficult to determine as the threshold pressure drop would have to be determined simultaneously with the exponent. This is where the mixed wet percolation problem enters. With the saturation (mixing ratio) being in the correct range, connected paths constituting pores between two opposite types of walls A and B would appear, where there would essentially be no capillary forces. Hence, the threshold pressure drop would be zero, and the exponent could be determined with good accuracy, $2.56\pm 0.05$, in this system [18].
It is the structure and transport properties of these $A-B$ paths, which form the percolation problem we focus on here. We refer to this variant of the percolation problem as mixed wet percolation.
@@ 2 The mixed-wet percolation (MWP) Model
Since the square lattice is a self-dual lattice, it is easy to model mixed wet percolation involving a primal square lattice and its self-dual lattice in $2$d. The ordinary site percolation is implemented on the primal lattice. Starting from an empty primal lattice, the sites are occupied randomly with a probability $p$ and the rest of the sites remain unoccupied with a probability $(1-p)$. The occupied and unoccupied sites represent $A$ and $B$ type grains, respectively. Then, we consider the links on the dual lattice. If a link of the dual lattice is either between two adjacent occupied sites or between two adjacent unoccupied sites on the primal lattice, no bond is placed at that link on the dual lattice. If, however, the two adjacent lattice sites on the primal lattice are of different types, i.e. occupied-unoccupied or vice versa, we place a bond on the link of the dual lattice with unit probability. It is demonstrated with a $2\times 2$ square lattice and two different grains in Fig. 2.
@@
Figure 2: The black circles represent $A$-type grains and the grey circles represent $B$-type grains. In (a), all grains are $B$-type. Hence, there will be no flow, and no links on the dual lattice will be occupied. Similarly, in (c), all the grains are $A$-type. It corresponds to no flow, and no links are occupied. In (b), two $A$-type and two $B$-type grains are placed diagonally opposite such that they are always adjacently different, and fluid can flow through all the links between them. Hence, they are occupied
@@
All the links on the dual lattice are then verified, and bonds are placed only if they are found to be between two adjacent sites of different types (occupied-unoccupied) on the primal lattice. Placing all such bonds, it is found that the occupied bonds on the dual lattice form perimeters of either the occupied site clusters or the unoccupied site clusters on the primal lattice. We call them perimeter bond clusters on the dual lattice. These bonds of mixed wet percolation represent the pores with zero capillary barriers, through which the two immiscible fluids are likely to flow in the two-component porous media in a capillary-dominated regime. We study the geometrical properties of these perimeter bond clusters. In Fig. 3(a) and (c), we present such perimeter bond clusters generated on the dual lattice of a $16\times 16$ square primal lattice for a site concentration $p=0.3$ and $p=0.7$, respectively.
@@
It can be seen that the perimeter bond clusters of mixed wet percolation are very different from random bond percolation (BP) clusters on a square lattice. The perimeter bond clusters are closed loops around the occupied (or unoccupied) sites without having a single open bond, in contrast to random bond percolation clusters. The smallest isolated cluster has four bonds, whereas, in random bond percolation, a single-bond cluster is possible. Unlike random bond percolation, all clusters have an even number of bonds. The perimeters around two clusters of occupied (or unoccupied) sites on the primal lattice are connected by a knot shown by red bullets in Fig. 3(a) and (c). A lattice site on the dual lattice will be called a knot if all four occupied bonds meet the site. It can be noticed that the perimeter bond clusters are very similar to the boundary of up-spin and down-spin domains in a spin-$1/2$ Ising system. At $p=0.3$, the perimeters are around the clusters of occupied sites on the primal lattice. This is because the density of occupied sites is less than that of the unoccupied sites. Symmetrically, at $p=0.7$, the perimeters are mostly around the clusters of unoccupied sites on the primal lattice, as their concentration is less than that of the occupied sites. Since a bond in the dual lattice appears between a pair of adjacent occupied and unoccupied sites on the primal lattice, the density of bonds $\rho$ in terms of the site density $p$ is given by
$\rho=2p(1-p)\;$ (1)
@@
where $\rho=B/(2L^{2})$, $B$ is the total number of occupied bonds in the square dual lattice of size $L\times L$ and the factor $2$ appears because a bond can also be generated by interchanging the positions of occupied-unoccupied sites, as shown in the inset of Fig. 3(b). In Fig. 3(b), $\rho$ is plotted against $p$, the site occupation probability. The blue line represents Eq. 1, and the black circles represent the measured bond density against $p$. A good agreement is observed. As $p$ increases from zero, the value of $\rho$ increases from zero, and it reaches a maximum value $\rho_{\rm max}=1/2$ at $p=1/2$ then decreases and goes to zero at $p=1$. The bond density is symmetric about $p=1/2$. As $\rho$ is a quadratic function of $p$, for a given $\rho$, there will be two values for $p$. Therefore, if there is a percolation threshold at which a perimeter bond cluster spans the dual lattice for the first time at $p=p_{c}<1/2$, then there must be another threshold at $p=1-p_{c}>1/2$ beyond which no perimeter bond cluster will span the dual lattice. To determine the critical thresholds ($p_{c}$), we will check whether a perimeter bond cluster spanned or percolated the dual lattice for a given $p$. The mixed wet percolation (or perimeter bond percolation) in the dual lattice will occur within these two percolation thresholds. First, we will be studying the geometrical properties of the mixed wet percolation clusters on the dual lattice around these thresholds. The results will be compared with that of ordinary percolation.
@@ 3 Scaling theory of geometrical quantities
$\phi(\rho)\sim\left|\Delta\rho\right|^{\pm q}$ (2)
where $q$ is an exponent, and $\pm$ corresponds to a singularity ($+$) or to a divergence (--). Now, we obtain $\Delta\rho$ in terms of $\Delta p=p-p_{c}$ as
$\displaystyle\Delta\rho$ $=2p(1-p)-2p_{c}(1-p_{c})=a\Delta p-b\left(\Delta p\right)^{2}$ (3)
where $a=2(1-2p_{c})$ and $b=2$. The above equation can also be obtained by a Taylor series expansion of $f(p)=2p(1-p)$ around $p=p_{c}$. Now, the bond cluster property $\phi(\rho)$ can be obtained in terms of $\Delta p$ as
$\phi(p)\sim\left|a\Delta p-b\left(\Delta p\right)^{2}\right|^{\pm q}=\left|\Delta p\right|^{\pm q}\left|\left(a-b\Delta p\right)\right|^{\pm q}$ (4)
@@
$\phi(p)\sim\left|\Delta p\right|^{\pm q}$ (6)
@@
As the site percolation scaling theory is well described [2, 3], we will provide the scaling relations for geometrical quantities related to mixed wet percolation clusters as a function of the control parameter $p$ around $p_{c}$ instead of the bond density $\rho$. A perimeter bond or mixed wet percolation cluster is characterized by the number of bonds $b$ present in that cluster and the cluster’s radius of gyration $R_{b}$. The geometrical quantities of mixed wet percolation clusters at a site concentration $p$ on the primal lattice of size $L\times L$ can be defined in terms of the cluster number density $n_{b}(p)$ [3, 2, 25]. The cluster number density $n_{b}(p)$ is defined as
$n_{b}(p)=\frac{N_{b}(p)}{2L^{2}}$ (7)
@@
$B_{\infty}(p)=\rho(p)-\sum_{b}{\vphantom{\sum}}^{\prime}bn_{b}(p)\sim(p-p_{c})^{\beta}$ (9)
@@
$\chi_{b}(p)=L^{d}\left[\langle B_{\infty}^{2}\rangle-\langle B_{\infty}\rangle^{2}\right]\sim|p-p_{c}|^{-\gamma}$ (10)
@@
$\xi_{b}^{2}(p)=2\sum_{b}{\vphantom{\sum}}^{\prime}R_{b}^{2}b^{2}n_{b}/\sum_{b}{\vphantom{\sum}}^{\prime}b^{2}n_{b}\sim|p-p_{c}|^{-2\nu}$ (11)
where $\nu=(\tau-1)/d\sigma$ is the correlation length exponent, $d$ is the space dimension and $R_{b}$ is the radius of gyration of a perimeter bond cluster of size $b$.
@@
$Q(p,L)=L^{q/\nu}\widetilde{Q}\left[(p-p_{c})L^{1/\nu}\right]$ (12)
where $q$ is a critical exponent, $\nu$ is the correlation length exponent and $\widetilde{Q}$ is the scaling function. Now, we provide FSS forms of some of the geometrical quantities we will use. The wrapping probability is the probability that a sample spans out of $N$ samples, and it is given by $W_{b}(p,L)=N_{\rm sp}/N$ where $N_{\rm sp}$ is the number of spanning samples out of $N$ samples. The FSS form of the wrapping probability $W_{b}(p,L)$ is given by
$W_{b}(p,L)=\widetilde{W}_{b}\left[(p-p_{c})L^{1/\nu}\right]$ (13)
@@
$B_{\infty}(p,L)=L^{-\beta/\nu}\widetilde{B}_{\infty}\left[(p-p_{c})L^{1/\nu}\right]$ (14)
where $\widetilde{B}_{\infty}$ is the scaling function. The FSS of fluctuation in order parameter $\chi_{b}$ is given by
$\chi_{b}(p,L)=L^{\gamma/\nu}\widetilde{\chi}_{b}\left[(p-p_{c})L^{1/\nu}\right]$ (15)
@@
The scaling behaviour of the geometrical quantities associated with the bond clusters of the RBMWP model should be studied in terms of $\Delta\omega=|\omega-\omega_{c}|$ where $\omega$ is bond occupation probability. The scaling relations of the geometrical quantities in RBMWP in terms of $\Delta\omega$ will be the same as described above.

### v2

@@ 1 Introduction
Percolation theory was first recognized as a critical phenomenon by Broadbent and Hammersley in 1957 broadbent1957percolation . Little did they know that their paper would give rise to literally thousands of papers in the ensuing years. It remains the foremost example of a non-thermal critical phenomenon stauffer2018introduction , complexity . The percolation problem in its primal form is easy to grasp. However, the percolation problem comes in many disguises. We may list some. Ambogaekar, Halperin and Langer demonstrated that the conductivity of strongly disordered conductors is related to a percolation threshold ambegaokar1971hopping . A similar line of reasoning for measuring the permeability of porous rocks is behind the well-known Katz and Thompson method of mercury injection katz1986quantitative . Slow injection of one fluid into a porous medium already containing another fluid, and the two fluids are immiscible, is well modeled by the invasion percolation model wilkinson1983invasion . A more recent version of the percolation problem that has generated considerable interest is explosive percolation achlioptas2009explosive , riordan2011explosive .
@@
We present here a new percolation model that is motivated by the flow of immiscible fluids in a mixed-wet porous medium. Percolation models have widely been used in the past and present to characterise porous media flow, especially in the capillary-dominated regime where the flow is governed by the disorder in the pores sahimi2011flow , blunt2017multiphase , feder2022physics , hs17 , ga19 . For example, the displacement of an immiscible fluid by another fluid inside a porous medium during capillary-dominated flow produces fingers due to capillary instabilities. These fingers are called capillary fingers, which were characterized by the invasion percolation model ltz88 , lz89 . The invasion percolation model has also been used to find the relative permeability and saturation profile near the percolation critical point for a static two-phase flow system in the capillary-dominated regime w86 . The percolation model we present here is motivated by the experimental studies irannezhad2023fluid , irannezhad2023characteristics , and the computational modeling fyhn2023effective of two-phase flow in an engineered mixed-wet porous medium that is made of a random mixture of two types grains with opposite wetting properties with respect to the fluids flowing in the medium. Imagine a random mixture of two types of grains, say $A$ and $B$, and two immiscible fluids, say fluid $1$ and fluid $2$. Suppose the grain type $A$ is less wetting to fluid $1$ compared to fluid $2$, i.e., the wetting angle $\theta_{2A}<\pi/2$, see Fig. 1. The grain type $B$, on the other hand, is more wetting to fluid $1$ compared to fluid $2$, i.e., $\theta_{1B}<\pi/2$. Suppose now we have an interface between fluid $1$ and $2$ moving in a pore with $A$-type walls. The capillary force between the two fluids will then point towards fluid $1$ as fluid $2$ is more wetting than fluid $1$, see Fig. 1 (left). If the walls are both of type $B$, the capillary force at the interface will point towards fluid $2$. However, if one wall is of type $A$ and the other of type $B$, we have the situation shown in Fig. 1 (right). In this case, the capillary forces will be much weaker than the two previous cases as the contribution from each wall essentially cancels.
@@
Geistlinger et al. geistlinger2021new did a micro-computed tomography ($\mu$-CT) study of a mixture of sand grains with contact angle either being $0^{\circ}$ or $100^{\circ}$ with respect to the fluid interfaces. They then injected one of the fluids into the ensuing porous medium saturated with the other fluid. They would recognize a structural percolation transition in the clusters of the defending fluid after the invasion process. Irannezahad et al. irannezhad2023fluid , irannezhad2023characteristics studied the same problem numerically but not in a percolation context. Fyhn et al. fyhn2023effective studied steady-state flow in mixed-wet porous media using a dynamic pore-network model sgv21 . By steady-state flow, we mean flow where the macroscopic averages of the various variables are constant or fluctuate around well-defined averages, and there are no macroscopic saturation gradients present erpelding2013history . Fyhn et al. fyhn2023effective recognized that the steady-state mixed wet problem is a percolation problem in disguise.
@@
In 2009, Tallakstad et al. tallakstad2009steady , tallakstad2009steadyb demonstrated that there is a flow regime under immiscible steady-state two-phase flow in porous media where the flow rate is proportional to the pressure gradient raised to a power close to two. This is due to interfaces held in place by capillary forces would be mobilized with increasing pressure drop. Sinha and Hansen sinha2012effective found that if pressure drop rather than flow rate was the control variable, there typically would be a threshold pressure drop for flow to occur. This would be caused by the immobilized interfaces blocking the flow. The presence of this threshold would make the determination of the exponent in the power law between flow rate and pressure drop difficult to determine as the threshold pressure drop would have to be determined simultaneously with the exponent. This is where the mixed wet percolation problem enters. With the saturation (mixing ratio) being in the correct range, connected paths constituting pores between two opposite types of walls $A$ and $B$ would appear, where there would essentially be no capillary forces. It is the structure of these paths between $A$ and $B$, which form the percolation problem we focus on here. We refer to this variant of the percolation problem as mixed wet percolation.
@@ 2 The mixed-wet percolation (MWP) Model
Since the square lattice is a self-dual lattice, it is easy to model mixed wet percolation involving a primal square lattice and its self-dual lattice in 22d. The ordinary site percolation is implemented on the primal lattice. Starting from an empty primal lattice, the sites are occupied randomly with a probability $p$ and the rest of the sites remain unoccupied with a probability (1p)(1-p). The occupied and unoccupied sites represent $A$ and $B$ type grains, respectively. Then, we consider the links on the dual lattice. If a link of the dual lattice is either between two adjacent occupied sites or between two adjacent unoccupied sites on the primal lattice, no bond is placed at that link on the dual lattice. If, however, the two adjacent lattice sites on the primal lattice are of different types, i.e. occupied-unoccupied or vice versa, we place a bond on the link of the dual lattice with unit probability. It is demonstrated with a $2\times 2$ square lattice and two different grains in Fig. 2.
@@
Figure 2: The black circles represent AA-type grains and the grey circles represent $B$-type grains. In (a), all grains are $B$-type. Hence, there will be no flow, and no links on the dual lattice will be occupied. Similarly, in (c), all the grains are AA-type. It corresponds to no flow, and no links are occupied. In (b), two AA-type and two $B$-type grains are placed diagonally opposite such that they are always adjacently different, and fluid can flow through all the links between them. Hence, they are occupied
@@
All the links on the dual lattice are then verified, and bonds are placed on a dual lattice link only if they are found to be between two adjacent sites of different types (occupied-unoccupied) on the primal lattice. Thus, a bond on the dual lattice is separating a pair of occupied-unoccupied sites on the primal lattice. Placing all such bonds, it is found that the occupied bonds on the dual lattice form perimeters of either the occupied site clusters or the unoccupied site clusters on the primal lattice. We call them perimeter bond clusters on the dual lattice. These bonds of mixed wet percolation represent the pores with zero capillary barriers, through which the two immiscible fluids are likely to flow in the two-component porous media in a capillary-dominated regime. We study the geometrical properties of these perimeter bond clusters. In Fig. 3(a) and (c), we present such perimeter bond clusters generated on the dual lattice of a $16\times 16$ square primal lattice for a site concentration $p=0.3$ and $p=0.7$, respectively.
@@
It can be seen that the perimeter bond clusters of MWP are very different from random bond percolation (BP) clusters on a square lattice. The perimeter bond clusters of MWP are closed loops around the occupied (or unoccupied) sites without having a single open bond, in contrast to random bond percolation clusters. The smallest isolated cluster has four bonds, whereas in random bond percolation, a single-bond cluster is possible. Unlike random bond percolation, all clusters have an even number of bonds. The perimeters around two clusters of occupied (or unoccupied) sites on the primal lattice are connected by a knot shown by red bullets in Fig. 3(a) and (c). A lattice site on the dual lattice will be called a knot if all four occupied bonds meet the site. It can be noticed that the perimeter bond clusters are very similar to the boundary of up-spin and down-spin domains in a spin-$1/2$ Ising system. At $p=0.3$, the perimeters are around the clusters of occupied sites on the primal lattice. This is because the density of occupied sites is less than that of the unoccupied sites. Symmetrically, at $p=0.7$, the perimeters are mostly around the clusters of unoccupied sites on the primal lattice, as their concentration is less than that of the occupied sites. Since a bond in the dual lattice appears between a pair of adjacent occupied and unoccupied sites on the primal lattice, the density of bonds $\rho$ in terms of the site density $p$ is given by
$\rho=2p(1-p)\;,$ (1)
@@
where $\rho=B/(2L^{2})$, $B$ is the total number of occupied bonds in the square dual lattice of size $L\times L$ and the factor $2$ appears because a bond can also be generated by interchanging the positions of occupied-unoccupied sites, as shown in the inset of Fig. 3(b). In Fig. 3(b), $\rho$ is plotted against $p$, the site occupation probability. The blue line represents Eq. 1, and the black circles represent the measured bond density against $p$. A good agreement is observed. As $p$ increases from zero, the value of $\rho$ increases from zero, and it reaches a maximum value $\rho_{\rm max}=1/2$ at $p=1/2$, then decreases and goes to zero at $p=1$. The bond density is symmetric about $p=1/2$. As $\rho$ is a quadratic function of $p$, for a given $\rho$, there will be two values for $p$. Therefore, if there is a percolation threshold at which a perimeter bond cluster spans the dual lattice for the first time at $p=p_{c}<1/2$, then there must be another threshold at $p=1-p_{c}>1/2$ beyond which no perimeter bond cluster will span the dual lattice. To determine the critical thresholds ($p_{c}$), we will check whether a perimeter bond cluster spanned or percolated the dual lattice for a given $p$. The mixed wet percolation (or perimeter bond percolation) in the dual lattice will occur within these two percolation thresholds. First, we will be studying the geometrical properties of the mixed wet percolation clusters on the dual lattice around these thresholds. The results will be compared with those of ordinary percolation.
@@
MWP should not be confused with the so-called $AB$ percolation wierman1989ab , wu2003ab in which $A$ and $B$ sites are connected by a bond on the primal lattice, and no dual lattice is involved. Moreover, $AB$ percolation does not exhibit a percolation transition on the square lattice. It is rather known as anti-percolation on the square lattice.
@@ 3 Scaling theory of geometrical quantities
$\phi(\rho)\sim\left|\Delta\rho\right|^{\pm q}\;,$ (2)
where $q$ is an exponent, and $\pm$ corresponds to a singularity (++) or to a divergence (-). Now, we obtain $\Delta\rho$ in terms of $\Delta p=p-p_{c}$ as
$\displaystyle\Delta\rho$ $=2p(1-p)-2p_{c}(1-p_{c})=a\Delta p-b\left(\Delta p\right)^{2}\;,$ (3)
where $a=2(1-2p_{c})$ and $b=2$. The above equation can also be obtained by a Taylor series expansion of f(p)=2p(1p)f(p)=2p(1-p) around $p=p_{c}$. Now, the bond cluster property $\phi(\rho)$ can be obtained in terms of $\Delta p$ as
$\phi(p)\sim\left|a\Delta p-b\left(\Delta p\right)^{2}\right|^{\pm q}=\left|\Delta p\right|^{\pm q}\left|\left(a-b\Delta p\right)\right|^{\pm q}\;,$ (4)
@@
$\phi(p)\sim\left|\Delta p\right|^{\pm q}\;,$ (6)
@@
As the site percolation scaling theory is well described stauffer2018introduction , complexity , we will provide the scaling relations for geometrical quantities related to mixed wet percolation clusters as a function of the control parameter $p$ around $p_{c}$ instead of the bond density $\rho$. A perimeter bond or mixed wet percolation cluster is characterized by the number of bonds $b$ present in that cluster and the cluster’s radius of gyration $R_{b}$. The geometrical quantities of mixed wet percolation clusters at a site concentration $p$ on the primal lattice of size $L\times L$ can be defined in terms of the cluster number density $n_{b}(p)$ complexity , stauffer2018introduction , herrmann1990modelization . The cluster number density $n_{b}(p)$ is defined as
$n_{b}(p)=\frac{N_{b}(p)}{2L^{2}}\;,$ (7)
@@
$B_{\infty}(p)=\rho(p)-\sum_{b}{\vphantom{\sum}}^{\prime}bn_{b}(p)\sim(p-p_{c})^{\beta}\;,$ (9)
@@
$\chi_{b}(p)=L^{d}\left[\langle B_{\infty}^{2}\rangle-\langle B_{\infty}\rangle^{2}\right]\sim|p-p_{c}|^{-\gamma}\;,$ (10)
@@
$\xi_{b}^{2}(p)=2\sum_{b}{\vphantom{\sum}}^{\prime}R_{b}^{2}b^{2}n_{b}\bigg/\sum_{b}{\vphantom{\sum}}^{\prime}b^{2}n_{b}\sim|p-p_{c}|^{-2\nu}\;,$ (11)
where $\nu=(\tau-1)/d\sigma$ is the correlation length exponent, $d$ is the space dimension, and $R_{b}$ is the radius of gyration of a perimeter bond cluster of size $b$.
@@
$Q(p,L)=L^{q/\nu}\widetilde{Q}\left[(p-p_{c})L^{1/\nu}\right]\;,$ (12)
where $q$ is a critical exponent, $\nu$ is the correlation length exponent and $\widetilde{Q}$ is the scaling function. Now, we provide FSS forms of some of the geometrical quantities we will use. The wrapping probability is the probability that a sample spans, and it is given by $W_{b}(p,L)=N_{\rm sp}/N$ where $N_{\rm sp}$ is the number of spanning samples out of $N$ samples. The FSS form of the wrapping probability $W_{b}(p,L)$ is given by
$W_{b}(p,L)=\widetilde{W}_{b}\left[(p-p_{c})L^{1/\nu}\right]\;,$ (13)
@@
$B_{\infty}(p,L)=L^{-\beta/\nu}\widetilde{B}_{\infty}\left[(p-p_{c})L^{1/\nu}\right]\;,$ (14)
where $\widetilde{B}_{\infty}$ is the scaling function. The FSS of fluctuation in the order parameter $\chi_{b}$ is given by
$\chi_{b}(p,L)=L^{\gamma/\nu}\widetilde{\chi}_{b}\left[(p-p_{c})L^{1/\nu}\right]\;,$ (15)
@@
The scaling behaviour of the geometrical quantities associated with the bond clusters of the RBMWP model should be studied in terms of $\Delta\omega=|\omega-\omega_{c}|$ where $\omega$ is the bond occupation probability. The scaling relations of the geometrical quantities in RBMWP in terms of $\Delta\omega$ will be the same as described above.